OpenStudy (anonymous):

need help solving for constants: my attempt (1+2s^2)/(s^2(s^2+3s+2)= A/(s^2)+B/(s+1)+C/(s+2) need help solving for constants: my attempt (1+2s^2)/(s^2(s^2+3s+2)= A/(s^2)+B/(s+1)+C/(s+2) @Mathematics

OpenStudy (anonymous):

I am pretty sure I did it wrong since it is s^2((s+1)(s+2))

OpenStudy (amistre64):

just gotta get another one in there for a single "s"

OpenStudy (amistre64):

you have 2 multiples of s; s^2 and s

OpenStudy (anonymous):

not sure I understand

OpenStudy (amistre64):

\[\frac{(1+2s^2)}{s^2(s^2+3s+2)}=\frac{A}{s^2}+\frac{B}{(s+1)}+\frac{C}{(s+2)}+\frac{D}{s} \]

OpenStudy (amistre64):

you have to account for all the possible decompositions

OpenStudy (amistre64):

1\s^2 decomps into 1/s and 1/s^2

OpenStudy (amistre64):

once thats determined; multiply both sides by the common denominator of the left

OpenStudy (amistre64):

\[\frac{(1+2s^2)}{s^2(s+1)(s+2)}=\frac{A}{s^2}+\frac{B}{(s+1)}+\frac{C}{(s+2)}+\frac{D}{s}\] \[s^2(s+1)(s+2)(\frac{(1+2s^2)}{s^2(s+1)(s+2)}=\frac{A}{s^2}+\frac{B}{(s+1)}+\frac{C}{(s+2)}+\frac{D}{s})\] \[1+2s^2=\]\[A(s+1)(s+2)+Bs^2(s+2)+Cs^2(s+1)+Ds(s+1)(s+2)\]

OpenStudy (anonymous):

Ohhh okay I see what to do now

OpenStudy (amistre64):

good ;)

OpenStudy (anonymous):

thanks for the help!

OpenStudy (amistre64):

yep, and good luck ;)